12.
The equation y = 10x – 70 describes the amount of money a class of students might earn from candy bar sales. What are the slope and y-intercept of this line?






Answer by venugopalramana(3286)


SEE THE FOLLOWING AND DO
------------------------------------
Find the slope and y-intercept.
10x – 7y = 7
PUT IT IN STANDARD FORM OF
Y=MX+C..WHERE..M=SLOPE AND C=Y INTERCEPT
SO WE GET
7Y=10X-7
Y=10X/7 - 1
SLOPE =10/7
Y INTERCEPT = - 1